the sum of two numbers is 84,and one of them is 12 more than the other.what are the two numbers?
Answer by erikaii.19(2)

a) let x+y=84
x=y+12



b) x+y=84
y+12+y=84
2y=84-12
2y=72
y=36


c) x=y+12
x=36+12
x=48
therefore the two nos. are 36 and 48. :)